OpenStudy (anonymous):

Name the property illustrated by each question. 5x · (4y + 3x) = 5x · (3x + 4y) I just need someone to jog my memory, I totally forgot how to do these.

OpenStudy (saifoo.khan):

cancel the both 5x on the sides.. u will be left with, 4y + 3x = 3x + 4y Add like terms. 4y - 4y = 3x - 3x 0 = 0 Answer.

OpenStudy (anonymous):

5x · (4y + 3x) = 5x · (3x + 4y) (4y + 3x) = (3x + 4y) 4y+3x=4y+3x 3x=3x, >>>>>0=0

OpenStudy (anonymous):

Commutativity of addition.